bob and fred play the following game. Bob rolls a single die. if the result is the number 1,2,3,4,5 bob must pay fred the number of dollars indicated by the number rolled. if a 6 is rolled fred must pay bob $57 dollars.
find freds expectation?

Answer by stanbon(75887)

bob and fred play the following game. Bob rolls a single die. if the result is the number 1,2,3,4,5 bob must pay fred the number of dollars indicated by the number rolled. if a 6 is rolled fred must pay bob $57 dollars.
find freds expectation?
---------------------------------
Random variable values: 1,2,3,4,5,-57
corresponding probabilities are each 1/6
-------------------------------------------
Expected "winnings": (1/6)(1+2+3+4+5-57) = (1/6)(-42) = -$7.00
===============================================================
Cheers,
Stan H.